# Magneto-optical properties of Group-IV–vacancy centers in diamond upon hydrostatic pressure

### v1

@@ Appendix B Hyperfine tensors under dynamic Jahn–Teller interaction
The hyperfine tensor elements can be evaluated by spinpolarized DFT whenever the $|e_{x}\rangle$ or $|e_{y}\rangle$ orbital is being occupied for defects exhibiting $D_{3d}$ symmetry. In simple words, the hyperfine tensor elements can be evaluated whenever $|e_{x}\rangle$ or $|e_{y}\rangle$ state is being selected. Therefore, the hyperfine interaction depicted in $\hat{H}_{\text{HF}}=\hat{S}A\hat{I}$ can be extended for orbital degrees of freedom as
$\hat{W}=$ $\overleftarrow{S}\overleftrightarrow{A}^{(xx)}\overrightarrow{I}|e_{x}\rangle\langle e_{x}|+\overleftarrow{S}\overleftrightarrow{A}^{(xy)}\overrightarrow{I}|e_{x}\rangle\langle e_{y}|$ (8)
@@
where $\overleftrightarrow{A}^{(xx)}$ and $\overleftrightarrow{A}^{(yy)}$ tensors can be determined by applying Eq. (2) upon spin densities of $|e_{x}\rangle$ and $|e_{y}\rangle$, respectively. However, $\overleftrightarrow{A}^{(xy)}$, $\overleftrightarrow{A}^{(yx)}$ offdiagonal terms cannot be computed directly. Therefore, we will utilize $D_{3d}$ symmetry constraints to deduce those non-trivial orbital flipping terms.
@@
$\hat{W}=\overleftarrow{S}\overleftrightarrow{A}\overrightarrow{I}\!+\!\overleftarrow{S}\overleftrightarrow{A}_{x}\overrightarrow{I}\hat{\sigma}_{z}\!+\!\overleftarrow{S}\overleftrightarrow{A}_{y}\overrightarrow{I}\hat{\sigma}_{x}\!+\!\underset{=0}{\underbrace{\overleftarrow{S}\overleftrightarrow{A}_{z}\overrightarrow{I}i\hat{\sigma}_{y}}}$
@@
Figure 7: Transformation of three equivalent carbon atoms. (a) and (b) represent the electronic wavefunction when the $|e_{x}\rangle$ and $|e_{y}\rangle$ are chosen, respectively. The red and blue circles represent the positive and negative values of $|e_{x}\rangle$ and $|e_{y}\rangle$ wavefunctions, respectively that we also quantify inside the […] parentheses. (c) and (d) depict the transformation laws for $\hat{C}_{3}$ and $\hat{C^{-1}}_{3}$ rotations, respectively. (e) depicts the $Z=$[111] oriented coordinates of the G4V defects. (f) depicts the Cartesian coordinates of diamond.
Next, we discuss how to calculate these hyperfine interactions in practice within spinpolarized DFT calculations. In the usual Born-Oppenheimer ab initio DFT calculations, the double degenerate $e_{g}$ orbital is occupied by a single electron in the ground state for G4V(--) defect. We analyze the system by assuming $e_{g(x)}$ is occupied by an electron in $D_{3d}$ symmetry configuration. Among the six equivalent immediate carbon neighbor atoms together with inversion symmetry, two pairs of three carbon atoms can be rotated to each other about the $C_{3}$ symmetry axis (see Fig. 7), e.g., $\mathrm{C_{1}}$, $\mathrm{C_{2}}$, $\mathrm{C_{3}}$ carbon atoms exhibit $\overleftrightarrow{A}^{(xx)}$, $\overleftrightarrow{B}^{(xx)}$, $\overleftrightarrow{C}^{(xx)}$ hyperfine tensors, respectively. In this example, we define the following hyperfine matrices for C1,
$\overleftrightarrow{A}^{(xx)}=\overleftrightarrow{A}^{(0^{\circ})}=\left\langle+1;0\right|\hat{W}\left|+1;0\right\rangle=\overleftrightarrow{A}+\overleftrightarrow{A}_{x}\text{.}$ (11)
@@
$\begin{split}\hat{C}_{3}\overleftrightarrow{B}^{(xx)}=\overleftrightarrow{A}^{(+120^{\circ})}=&\biggl{\langle}-\frac{1}{2};-\frac{\sqrt{3}}{2}\biggr{|}\hat{W}\biggl{|}-\frac{1}{2};-\frac{\sqrt{3}}{2}\biggr{\rangle}\\ &=\overleftrightarrow{A}-\frac{1}{2}\overleftrightarrow{A}_{x}+\frac{\sqrt{3}}{2}\overleftrightarrow{A}_{y}\text{,}\end{split}$ (12a)
$\begin{split}\hat{C}_{3}^{-1}\overleftrightarrow{C}^{(xx)}=\overleftrightarrow{A}^{(-120^{\circ})}=&\biggl{\langle}-\frac{1}{2};+\frac{\sqrt{3}}{2}\biggr{|}\hat{W}\biggl{|}-\frac{1}{2};+\frac{\sqrt{3}}{2}\biggr{\rangle}\\ &=\overleftrightarrow{A}-\frac{1}{2}\overleftrightarrow{A}_{x}-\frac{\sqrt{3}}{2}\overleftrightarrow{A}_{y}\text{.}\end{split}$ (12b)
@@
where we also included the effect of $q=(1+p)/2$ vibronic reduction factor Bersuker (2006); Bersuker and Polinger (2012) that partially quenches the strength of $\{\hat{\sigma}_{z},\hat{\sigma}_{x}\}$ orbital operators transforming as $E$ representation of $D_{3d}$. The $p$ reduction factor reduces the orbital operators transforming as $A_{2}$ such as $\hat{L}_{z}=\hat{\sigma}_{y}$ of Eq. (1). We note that the $\overleftrightarrow{A}$ tensor can be evaluated by putting two half electrons on $|e_{x}\rangle$ and $|e_{y}\rangle$ orbitals which method was already used in Ref. Harris et al., 2023. Finally, $\overleftrightarrow{A}_{x}$ and $\overleftrightarrow{A}_{y}$ are not independent for the central atom in the inversion center. Therefore, the dopant atom’s hyperfine tensor can be represented by four hyperfine parameters, $A_{\parallel,\perp,1,2}$ that can be given Thiering and Gali (2024) as
$\hat{W}=$ $A_{\parallel}\hat{S}_{z}\hat{I}_{z}+A_{\perp}\left(\hat{S}_{x}\hat{I}_{x}+\hat{S}_{y}\hat{I}_{y}\right)+$ (14)
$A_{1}\left[(\hat{S}_{x}\hat{I}_{z}+\hat{S}_{z}\hat{I}_{x})\hat{\sigma}_{z}+(\hat{S}_{y}\hat{I}_{z}+\hat{S}_{y}\hat{I}_{x})\hat{\sigma}_{x}\right]+$
$A_{2}\left[(\hat{S}_{y}\hat{I}_{y}-\hat{S}_{x}\hat{I}_{x})\hat{\sigma}_{z}+(\hat{S}_{x}\hat{I}_{y}+\hat{S}_{y}\hat{I}_{x})\hat{\sigma}_{x}\right]\text{.}$
This equation can be rewritten with the definition $\sigma_{\pm}=\langle e_{\mp}|e_{\pm}\rangle$ and the usual ladder operators which then reads as
$\hat{W}=$ $A_{\parallel}\hat{S}_{z}\hat{I}_{z}+\frac{1}{2}A_{\perp}\left(\hat{S}_{+}\hat{I}_{-}+\hat{S}_{-}\hat{I}_{+}\right)+$ (15)
$A_{1}\left[(\hat{S}_{+}\hat{I}_{z}+\hat{S}_{z}\hat{I}_{+})\hat{\sigma}_{-}+(\hat{S}_{z}\hat{I}_{-}+\hat{S}_{-}\hat{I}_{z})\hat{\sigma}_{+}\right]+$
$A_{2}\left[(\hat{S}_{-}\hat{I}_{-}\hat{\sigma}_{-}+\hat{S}_{+}\hat{I}_{+})\hat{\sigma}_{+}\right]\text{.}$
@@
Table 6: The raw data of hyperfine tensors in both $D_{3d}$ and $C_{2h}$ symmetries for G4V(--) defects at zero pressure as obtained in HSE06 calculations in the electronic excited state. $A_{xx}$, $A_{yy}$ and $A_{zz}$ are the symmetric hyperfine tensor elements. $A_{xy}$, $A_{xz}$ and $A_{yz}$ parameters are the asymmetric hyperfine tensor elements. The unitless reduction factor $q$ is given for each defect associated with the Jahn–Teller effect. Atoms are labeled for low $C_{2h}$ symmetry configuration in Fig. 6(j) and the spin density of high $D_{3d}$ symmetry configuration is illustrated in Fig. 6.

### v2

@@ Appendix D Hyperfine interaction for 13C neighbors
In the previous Appendix C we discussed thoroughly the case of the central atom. However, we need a more generic form for 13C because they are not sitting on the central high symmetry point that of $D_{3d}$. As before, hyperfine tensor elements can be evaluated by spin-polarized DFT whenever the $|e_{x}\rangle$ or $|e_{y}\rangle$ orbital is being occupied for defects exhibiting $D_{3d}$ symmetry. In simple words, hyperfine tensor elements can be evaluated whenever the $|e_{x}\rangle$ or $|e_{y}\rangle$ state is selected. Therefore, the hyperfine interaction depicted can be extended for orbital degrees of freedom as
$\hat{H}_{\text{HF}}^{\mathrm{{}^{13}C}}\!=$ $\overleftarrow{S}\overleftrightarrow{A}^{(xx)}\overrightarrow{I}|e_{x}\rangle\langle e_{x}|+\overleftarrow{S}\overleftrightarrow{A}^{(xy)}\overrightarrow{I}|e_{x}\rangle\langle e_{y}|$ (30)
@@
where $\overleftrightarrow{A}^{(xx)}\bigr{|}_{ij}=\mathcal{A}_{ij}^{x}$ and $\overleftrightarrow{A}^{(yy)}\bigr{|}_{ij}=\mathcal{A}_{ij}^{y}$ tensors can be determined by applying Eq. (3) upon spin densities of $|e_{x}\rangle$ and $|e_{y}\rangle$, respectively. However, $\overleftrightarrow{A}^{(xy)}$, $\overleftrightarrow{A}^{(yx)}$ off-diagonal terms cannot be computed directly. We also cannot act similarly to what we did in Eq. (14) because 13C ions are not sitting in the central position.
@@
$\hat{H}_{\text{HF}}^{\mathrm{{}^{13}C}}\!=\overleftarrow{S}\overleftrightarrow{A}\overrightarrow{I}\!+\!\overleftarrow{S}\overleftrightarrow{A}_{x}\overrightarrow{I}\hat{\sigma}_{z}\!+\!\overleftarrow{S}\overleftrightarrow{A}_{y}\overrightarrow{I}\hat{\sigma}_{x}\!+\!\underset{=0}{\underbrace{\overleftarrow{S}\overleftrightarrow{A}_{z}\overrightarrow{I}i\hat{\sigma}_{y}}}$
@@
Figure 8: Transformation of three equivalent carbon atoms. (a) and (b) represent the electronic wavefunctions when the $|e_{x}\rangle$ and $|e_{y}\rangle$ are chosen, respectively. The red and blue circles represent the positive and negative values of $|e_{x}\rangle$ and $|e_{y}\rangle$ wavefunctions, respectively that we also quantify inside the […] parentheses. (c) and (d) depict the transformation laws for $\hat{C}_{3}$ and $\hat{C^{-1}}_{3}$ rotations, respectively. (e) depicts the $z=$[111] oriented coordinates of the G4V defects. (f) depicts the Cartesian coordinates of diamond. (g) Depiction of single particle orbitals of G4V centers. Green/purple color depicts positive/negative wavefunction phase. We also show the subduced representation labels from $D_{3v}$ to $C_{3v}$.
Next, we discuss how to calculate these hyperfine interactions in practice within spin-polarized DFT calculations. In the usual Born-Oppenheimer ab initio DFT calculations, the double degenerate $|e_{\{g,u\}}\rangle$ orbital is occupied by a single electron in the ground state for G4V(--) defect. We analyze the system by assuming that $|e_{{\{g,u\}}(x)}\rangle$ is occupied by an electron in the symmetry configuration $D_{3d}$. Among the six equivalent immediate carbon neighbor atoms together with inversion symmetry, two pairs of three carbon atoms can be rotated to each other about the $C_{3}$ symmetry axis (see Fig. 8), e.g., $\mathrm{C_{1}}$, $\mathrm{C_{2}}$, $\mathrm{C_{3}}$ carbon atoms exhibit $\overleftrightarrow{A}^{(xx)}$, $\overleftrightarrow{B}^{(xx)}$, $\overleftrightarrow{C}^{(xx)}$ hyperfine tensors, respectively. In this example, we define the following hyperfine matrices for C1,
$\overleftrightarrow{A}^{(xx)}=\overleftrightarrow{A}^{(0^{\circ})}=\left\langle+1;0\right|\hat{H}_{\text{HF}}^{\mathrm{{}^{13}C}}\left|+1;0\right\rangle=\overleftrightarrow{A}+\overleftrightarrow{A}_{x}\text{,}$ (33)
where $\left\langle+1;0\right|$ and $\left|+1;0\right\rangle$ indicate being fully in $|e_{x}\rangle$ orbital than $|e_{y}\rangle$.
@@
$\begin{split}\hat{C}_{3}\overleftrightarrow{B}^{(xx)}=\overleftrightarrow{A}^{(+120^{\circ})}=&\biggl{\langle}-\frac{1}{2};-\frac{\sqrt{3}}{2}\biggr{|}\hat{H}_{\text{HF}}^{\mathrm{{}^{13}C}}\biggl{|}-\frac{1}{2};-\frac{\sqrt{3}}{2}\biggr{\rangle}\\ &=\overleftrightarrow{A}-\frac{1}{2}\overleftrightarrow{A}_{x}+\frac{\sqrt{3}}{2}\overleftrightarrow{A}_{y}\text{,}\end{split}$ (34a)
$\begin{split}\hat{C}_{3}^{-1}\overleftrightarrow{C}^{(xx)}=\overleftrightarrow{A}^{(-120^{\circ})}=&\biggl{\langle}-\frac{1}{2};+\frac{\sqrt{3}}{2}\biggr{|}\hat{H}_{\text{HF}}^{\mathrm{{}^{13}C}}\biggl{|}-\frac{1}{2};+\frac{\sqrt{3}}{2}\biggr{\rangle}\\ &=\overleftrightarrow{A}-\frac{1}{2}\overleftrightarrow{A}_{x}-\frac{\sqrt{3}}{2}\overleftrightarrow{A}_{y}\text{.}\end{split}$ (34b)
@@
where we also included the effect of $q=(1+p)/2$ vibronic reduction factor [55, 56] that partially quenches the strength of $\{\hat{\sigma}_{z},\hat{\sigma}_{x}\}$ orbital operators transforming as $E$ representation of $D_{3d}$. The $p$ reduction factor reduces the orbital operators transforming as $A_{2}$ such as $\hat{L}_{z}=\hat{\sigma}_{y}$ of Eq. (1). We note that the $\overleftrightarrow{A}$ tensor can be evaluated by putting two half electrons on $|e_{x}\rangle$ and $|e_{y}\rangle$ orbitals which method was already used in Ref. 35.
@@
Table 11: The raw data of hyperfine tensors in both $D_{3d}$ and $C_{2h}$ symmetries for G4V(--) defects at zero pressure as obtained in HSE06 calculations in the electronic excited state. $A_{xx}$, $A_{yy}$ and $A_{zz}$ are the symmetric hyperfine tensor elements. $A_{xy}$, $A_{xz}$ and $A_{yz}$ parameters are the asymmetric hyperfine tensor elements. The reduction factor $q$ is given for each defect associated with the Jahn–Teller effect. Atoms are labeled for low $C_{2h}$ symmetry configuration in Fig. 7(j) and the spin density of high $D_{3d}$ symmetry configuration is illustrated in Fig. 7.